OpenStudy (anonymous):

The center of the ellipse is at: (-3, -6) (3, 6) (3, 4)

OpenStudy (anonymous):

\[\frac{ (x-3)^2 }{ 9 }+\frac{ (y-6)^2 }{ 16 }=1\]

OpenStudy (the_joker<3):

WHAT DO YOU THINK IT IS? >:}

OpenStudy (anonymous):

I have absolutely NO idea. I don't understand ellipses..lol

OpenStudy (anonymous):

Maybe it's A? Idk

OpenStudy (the_joker<3):

I THINK YOUR RIGHT >:}}}

OpenStudy (anonymous):

Anybody else have opinions? Lol

OpenStudy (the_joker<3):

ACTUALITY ID GO WITH B!

Nnesha (nnesha):

(h,k) is the center of the equation

OpenStudy (anonymous):

Idk what (h,k) is.

Nnesha (nnesha):

\[\huge\rm \frac{ (x-h)^2 }{ a^2}+\frac{ (y-k)^2 }{ b^2 }=1\] standard form of ellipse where (h,k) is the center

Nnesha (nnesha):

so what is h and k in the equation ?

OpenStudy (anonymous):

Still don't know.

OpenStudy (anonymous):

the answer will be B

Nnesha (nnesha):

the number with x in the parentheses is suppose to be h and number with y in the parentheses is k

Nnesha (nnesha):

sigh.

OpenStudy (anonymous):

What?

OpenStudy (the_joker<3):

THE ANSWER IS B THATS WHAT I SAID

Nnesha (nnesha):

.

OpenStudy (anonymous):

What's that supposed to mean? @Nnesha

OpenStudy (the_joker<3):

COULD I GET A MEDAL ? PLEAS